OpenStudy (anonymous):

What is the derivative of x^(2/3)(x+2)

OpenStudy (anonymous):

\[X ^{2/3}(x+2)\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@phi @robtobey

OpenStudy (phi):

did you try, using the product rule ?

OpenStudy (anonymous):

No! I forgot!! LOL

OpenStudy (anonymous):

Thanks!!!!!

OpenStudy (phi):

or, distribute the x^(2/3) and find the derivative of \[ x^\frac{5}{3} + 2x^\frac{2}{3} \]

OpenStudy (anonymous):

Ok, how do you do that? I made a mistake somewhere when I was doing that

OpenStudy (anonymous):

I brought the exponents down and subtracted the powers by 1

OpenStudy (phi):

write x^1 as x^(3/3) x^(2/3) * x^(3/3) = x^(2/3 + 3/3) i.e. add exponents

OpenStudy (anonymous):

Where'd you get x^1??

OpenStudy (phi):

adding exponents is easy to remember if you remember x^1 * x^2 (for example) is x*x*x= x^3

OpenStudy (anonymous):

Oh, yeah

OpenStudy (phi):

x^(2/3)(x+2)

OpenStudy (anonymous):

Ok

OpenStudy (phi):

distribute the x^2/3 x^2/3 * x + 2 x^(2/3)

OpenStudy (anonymous):

So x^5/3+4/3x^-1/3

OpenStudy (anonymous):

Oh whoops

OpenStudy (anonymous):

Its 5/3x^2/3+4/3x^-1/3

OpenStudy (phi):

yes, that is one way to write it.

OpenStudy (anonymous):

Ok, now how do I simplify that

OpenStudy (anonymous):

I need to find the critical points

OpenStudy (phi):

set it equal to 0

OpenStudy (anonymous):

Well to do that, I would like to simplify it further. Thats where I was making my mistakes

OpenStudy (anonymous):

I got (5x^2/3+4)/(3cuberoot(x)), but there is not supposed to be an exponent in the numerator

OpenStudy (phi):

\[ \frac{5}{3} x^\frac{2}{3} + \frac{4}{3} x^{-\frac{1}{3} } = 0 \] multiply both sides by 3 \[ 5 x^\frac{2}{3}+ 4 x^{-\frac{1}{3} } = 0 \] multiply both sides by x^(1/3)

OpenStudy (anonymous):

What I meant was to convert the -1/3 to 1/cube root of x

OpenStudy (anonymous):

Its easier for me that way because I can set the numerator equal to 0 to find the places where the derivative is 0, and so I can easily set the denominator equal to 0 to find the places where the derivative is undefined

OpenStudy (phi):

I would try the way I am suggesting

OpenStudy (anonymous):

Can we try it my way? I don't want to get confused and I like to be able to tell where the derivatives 0s are and where its undefined

OpenStudy (phi):

when working with fractional exponents, it is easiest to keep them as exponents. In other words, do not write it as cube root. starting with the derivative set equal to 0: \[ \frac{5}{3} x^\frac{2}{3} + \frac{4}{3} x^{-\frac{1}{3} } = 0 \] what did you do ? Please go in small steps so I can follow.

OpenStudy (phi):

btw, the only way to get a "divide by 0" is when x=0.

OpenStudy (anonymous):

I did \[((5x ^{2/3})/3)+4/(3\sqrt[3]{x})\]

OpenStudy (phi):

set equal to 0

OpenStudy (anonymous):

Actually, I'm not sure how I got that

OpenStudy (phi):

but you really should write it as x^(1/3) not cube root.

OpenStudy (anonymous):

Why? Shouldn't I combine like terms first

OpenStudy (phi):

now what ?

OpenStudy (anonymous):

I want to just have a numerator over a denominator, not two different fractions

OpenStudy (phi):

you need a common denominator

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

5x/3cube root of x + 4/3 cube root of x

OpenStudy (anonymous):

Ooooooh

OpenStudy (anonymous):

I don't think I found a common denominator. I think i just saw the two fractions (4/3 and 5/3) and said that 3 will be the denominator

OpenStudy (phi):

as long as x is not 0, you can ignore the bottom and solve for the top=0

OpenStudy (anonymous):

Yeah I get it now! Thanks a lot!!! These careless mistakes kill!

OpenStudy (phi):

fyi, if you did this \[ 5 x^\frac{2}{3}+ 4 x^{-\frac{1}{3} } = 0 \\ 5 x^\frac{2}{3}x^\frac{1}{3}+ 4 x^{-\frac{1}{3} }x^\frac{1}{3} = 0 \\5x + 4 =0 \\ x = - \frac{4}{5} \]